OpenStudy (anonymous):

Last question f(x) = x + 2 and g(x) = x^2 – 6x + 3. Find g(f(x))

OpenStudy (anonymous):

is it x^2-5x+5 ??

OpenStudy (mr.math):

g(f(x)) means that you plug f(x) in every x you have in g(x). \[g(f(x))=(x+2)^2-6(x+2)+3=x^2+4x+4-6x-12+3\] \[=x^2-2x-5\]

OpenStudy (anonymous):

x+ (-6) =-5

OpenStudy (hoblos):

it is x^2 -2x -5

OpenStudy (anonymous):

thank you guyz for the help